# A Survey of Passive Sensing in the Workplace

### v2

@@ Abstract
As emerging technologies increasingly integrate into all facets of our lives, the workplace stands at the forefront of potential transformative changes. A notable development in this realm is the advent of passive sensing technology, designed to enhance both cognitive and physical capabilities by monitoring human behavior. This paper reviews current research on the application of passive sensing technology in the workplace, focusing on its impact on employee wellbeing and productivity. Additionally, we explore unresolved issues and outline prospective pathways for the incorporation of passive sensing in future workplaces.
@@
**footnotetext: These authors contributed equally to this work.
I Introduction
With the growing adoption of technologies in the workplace, research on the “Future of Work” is intensively exploring ways to intelligently redesign workspaces. The goal is to enhance the talents of the workforce, both those recorded on and off the balance sheet. Among the rapidly growing technologies, passive sensors stand out for their ability to seamlessly integrate into work environments. These sensors collect data in situ without necessitating active engagement from individuals, facilitating a non-intrusive and pervasive means of understanding and supporting workers [1]. Through the passive collection of data, these technologies offer nuanced insights into worker behavior, allowing for empirical, contextually grounded analyses. Researchers have drawn links between passive sensing data and various behavioral markers, including mental wellbeing [2], personality traits [3], and productivity levels [4]. In the realm of the Future of Work, such technologies are invaluable for examining workers’ experiences both within and beyond the workplace over extended periods and at a broad scale. This is particularly relevant when traditional methods of evaluating worker performance and satisfaction—relying on self-reports or assessments from colleagues and supervisors—may fall short. The objective data collected from passive sensing technologies signals the emergence of innovative, hands-off approaches to assessing work dynamics.
We present a survey of existing research on the use of passive sensing technologies in the workplace to assess and promote wellbeing and productivity of the workforce. In this work, we consider “workplace” as the setting or place of employment where individuals perform tasks for their employer without regards to whether the space is private or shared. The results point out open problems and possible future directions in the field and suggest challenges to furthering the Future of Work using passive sensing technologies.
@@ II SURVEY METHODOLOGY
To ensure the relevance and rigor of our research, we conducted searches for terms like “passive sensing & future of work”, “wearables & worker productivity”, and “wearables & employee health”, leading us to review over 25,000 articles on platforms like Google Scholar. Our selection criteria favored recent papers published since 2015, including some seminal earlier works, and prioritized publications with a Google h5 index of 20 or higher. Furthermore, we imposed specific conditions to refine our focus. Studies had to involve the actual workforce engaged in day-to-day job activities, excluding simulations with non-working groups, such as students in office environments. The technologies considered had to be genuinely passive—unobtrusive, portable, and comfortably integrated into daily routines without requiring active interaction. The technologies should blend with the workers’ routine, rather than the workers changing their everyday routine to satisfy the study requirements. Therefore, research utilizing intrusive devices like EEG headsets, chest straps, gaze trackers, cameras and other cumbersome scientific devices was excluded in favor of those using more discreet and workplace-appropriate technologies. We focus on devices that we expect to have a higher acceptance in the office environment. This stringent selection process allowed us to accurately portray the state-of-the-art developments, progress, and hurdles in this field, providing a foundation to stimulate and guide further research on passive sensing technologies in the workplace.
III PASSIVE SENSING
As ubiquitous devices become more embedded in our lives, they offer an unprecedented ability to passively capture our daily behavior at a high resolution via multiple sensors. Researchers have been leveraging passive sensing techniques to model users’ behavior and their environment from various perspectives. Before the emergence of smartphones and wearable devices, sensors built in cell phones (such as GPS and Bluetooth) were employed to passively sense users’ context. RealityMining [5] uses location signals and Bluetooth log data to recognize social patterns in daily user activity, identify socially significant locations and model organizational rhythms. CenceMe [6] uses a wider range of sensors – including embedded cameras, microphones, accelerometer, GPS, temperature, light, humidity, magnetometer and button clicks – to detect users’ activities and habits and injected these into popular social networking applications to promote social interaction. In addition to capturing personal behavior, passive sensing devices at a large scale enabled people-centric urban sensing. MetroSense [7] combines mobile sensors (carried by people) and static sensors (installed among a city) to establish a large-scale passive sensing network that could support a wide variety of applications such as enterprises and hospitals. Recently, the advancement of mobile and wearable technology has further enhanced the capability of ubiquitous devices through longitudinal passive sensing. For example, StudentLife [2] uses smartphones to track college students’ daily behaviors for more than two months to assess their mental health, academic performance and behavioral trends.
Passive sensing is increasingly involved in various aspects of our daily lives. Within the workplace, it has been used to monitor physiological factors of workers, promote work safety, enhance efficiency among other things [8]. Recently, the Tesserae [9] project involved over 700 information workers to investigate how such technologies can be leveraged to measure workplace performance such as organizational citizenship behavior, as well as psychological traits and physical characteristics. In the rest of the article, we focus on applications of passive sensing technologies in the workplace to assess wellbeing and productivity of the workforce as the transition to the Future of Work takes place.
@@ III PASSIVE SENSING
TABLE I: Studied behavior and sensors used. The following table lists all the behaviors studied in the papers we surveyed and the corresponding sensor data used.
@@
Skin conductance, skin temperature, acceleration [13]; workplace behavioral logs/ computer activity [14, 15];
@@
EDA, blood volume pulse (BVP), acceleration, phone usage [40]; Computer activity [28, 30]; Speech, calendar [30];
@@
IV Understanding Workplace Dynamics: Insights from Passive Sensing on Wellbeing and Productivity
@@
In this section, we examine the pivotal role of passive sensing technologies in deepening our understanding of workplace wellbeing and productivity. Through an examination of recent research studies, we highlight how these innovative tools offer important insights into the complex interplay between work environments and employee behavior. This exploration sheds light on the potential of passive sensing to optimize workplace dynamics and underscores its significance in fostering a more productive and health-conscious work culture.
IV-A WELLBEING
@@
Over the years, there has been an increase in studies that explore the sensing capabilities of smartphones and wearables to assess as well as improve worker’s health and wellbeing. Among the papers selected, mental health issues such as stress, anxiety, affects are important dimensions studied. Most of these studies ask participants to self-report their health and wellbeing using validated instruments. While some studies use one-off survey scores, others employ experience sampling to study worker’s wellbeing in a longitudinal fashion. We list the key studies in Table I.
@@
Assessing workers’ stress and/or anxiety is perhaps the most studied topic about worker’s wellbeing. As stress and anxiety are known to have an impact on heart rate (HR), most of the studies rely on HR-based signals obtained from wrist-worn wearables that employ photoplethysmography (PPG) [45] sensors. For instance, Feng et al. [10] show that PPG based HR and step counts from a Fitbit can be used to classify the anxiety level of nursing professionals. The authors introduce a pipeline for discovering behavioral consistency features, the inclusion of which leads to an overall improvement in the predictive model, with an accuracy of about 58%. For this work, the authors use the median-splitted score of self-reported State-Trait Anxiety Inventory (STAI) [46] as the ground truth. Similarly, in a study of 10 construction workers, Jebelli et al. [11] use a wrist wearable capable of not only capturing the participant’s HR through PPG signals, but also their electrodermal activity (EDA) and peripheral skin temperature (ST). The study uses Salivary cortisol as the ground truth to assess the baseline stress levels of the workers, obtaining an accuracy of about 73% when distinguishing between low, medium and high stress level of the workers. Similarly, Mark et al. [29] discovered in their research on information workers that there was a direct correlation between the daily time spent on emails (captured via activity logging software) and increased levels of reported stress. In a bid to craft a stress-reduction intervention for the workplace, Howe et al. [12] conducted a four-week longitudinal study involving 86 participants to explore how the type and timing of interventions influence their usage, effectiveness in reducing stress, and user preferences. The authors utilized various signals including email, calendar entries, time of day, facial images, and heart rate data in their analysis. The authors conclude that digital micro-interventions successfully alleviate short-term stress in the workplace, advocating for their immediate incorporation into work settings to yield positive effects.
Another health-related issue that is of frequent interest is sleep. In one study of 138 individuals working as full-time nursing professionals, Feng et al. [19] use PPG based HR from Fitbit, sleep duration, sleep efficiency, REM Sleep duration features to gauge sleep quality of the participants. The study uses the Pittsburgh Sleep Quality Index (PSQI) [47] survey responses from participants as the ground truth. A score above 7 is taken as one class (potential sleep disorders) and a lower score is associated with the remaining class. The authors propose an optimization problem to identify the key meaningful motifs. Motif features obtained using the proposed pipeline results in higher performance while predicting sleep quality – while Fitbit summary results in 61.90% F1 score, the motif extracted from the proposed pipeline results in 77.47% F1 score. As wearable-based sleep detection is increasingly being used and relied on, one research work investigates its accuracy while tracking 700 information workers throughout a year using wearables and phones [20]. Martinez et al. report that wearables can overestimate sleep and they propose fusing phone usage activity (i.e., whether the phone is being used) with wearable data to mitigate this. Using the proposed approach, authors generate better models of self-reported sleep for the information workers than using either stream alone.
Worker’s affect, emotions and mood have also been studied in this domain. For example, Bin Morshed et al. [21] explore mood instability of 603 information workers and find that the sleep and activity duration obtained from a Garmin wearable are negatively correlated with mood instability scores. The authors compute mood instability in the study using self-reported positive and negative affect schedule (PANAS) [48]. Umematsu et al. [13] use wrist wearables that capture skin conductance, skin temperature and acceleration from 39 workers in a Japanese company over a 30-day period. Stress, mood and health scores are collected every morning, using self-reported scores from 0 to 100. The authors then try forecasting the scores for the next day using the previous day’s physiological data. They obtain a mean absolute error of 13.47, 14.09 and 18.51 while predicting stress, mood and health score respectively. For fine-grained mood classification, Zenonos et al.[26] proposed a framework that classifies eight different types of moods into five categories using sensor data such as heart rate, pulse rate, PPG, ECG, skin temperature, etc. Mark et al. [22] studied how workers’ moods were connected to the amount of time they spent on emails. They found that email usage had a negative correlation with mood balance: the more time someone spent on emails, the worse their mood became in comparison to their positive feelings.
@@
In another study of 50 hospital workers which also uses PANAS, Nadarajan et al. [23] find that speech activity can explain some variance in predicting positive affect measure. Employees wear a specifically designed audio badge during their work-shift hours. The authors extract several features from the audio to identify foreground speech. They then use a linear mixed effects model to estimate positive and negative affect from foreground activation (i.e., the percentage of recording time that foreground speech is present). Similarly, in another study, Robles-Granda et al. [16] utilize multiple sensing modalities to assess anxiety, sleep and affect of 757 information workers. Sensing modalities include wearable, phone application, Bluetooth beacons and social media. Models trained on the fusion of all the features from different sensing modalities leads to up to 13.9% improvement in the symmetric mean absolute percentage error (SMAPE) when predicting affect, anxiety and sleep quality scores.
In yet another study, Booth et al. employ multiple sensing modalities to study stress, anxiety and affect of hospital workers [18]. The ground truth for stress and anxiety are obtained as self-reports on a 5-point Likert scale, whereas PANAS is used for assessing affect. The work investigates how patterns in movement data in the workplace can relate with mental wellness. The hospital staffs carry a Bluetooth tag with them which communicates with several Bluetooth hubs placed in different rooms of the hospital. The nearest hub which has the highest RSSI is taken as the location estimate. For each consecutive data in which the participant is in the same location, authors record the entry time, the location and the duration spent there. The authors extract motifs from each time series of locations where participants linger and cluster these motifs to capture temporal relationships between rooms. Authors obtain 58% accuracy while classifying the mental wellness of the participants (between a binary label with one value representing high stress, negative affect, and anxiety, and the other value representing low stress and positive affect) using the combination of extracted motif features with Fitbit data (step count, heart rate and sleep) and smart shirt data (heart rate, breathing, acceleration etc.).
@@
Other health and wellbeing related topics that have been studied using passive sensing among workers include focus and awakeness. Soto et al. utilize biometric data from an arm-wear (viz., physical activity, HR, skin response, skin temperature and respiration) to estimate worker’s stress, focus and awakeness [25]. The study ran for 8 weeks with 14 knowledge workers working at a large power and automation company. The ground truth for awakeness and focus along with stress were obtained as self reports from the participants. The authors report that personalized models outperformed generic model, predicting baselines ranging from 3% improvement to up to 52% improvement in the precision, recall and accuracy score. There are also several other works that fuse passive sensing data with some other modality to assess wellbeing. For instance, Saha and Grover et al. [49] propose contextualizing on offline behaviors as obtained from passive sensors to make models better adapted to the social media signals. Authors first cluster participants based on the passive sensing data obtained from their wearable, Bluetooth beacons and smartphone application. Thereafter, they use social media derived features to predict different constructs within each cluster, obtaining an improvement on the baseline generic model by up to 5.43% in predicting self-reported anxiety, affect and sleep quality of information workers.
@@
IV-B PRODUCTIVITY
Driven by the early success in the assessment of wellbeing in organizations, researchers have turned their focus to the use of passive sensing in the assessment of workplace performance. By tracking behaviors, recent work has attempted to characterize workplace performance with the long-term goals of identifying the behaviors that support productivity as well as generating interventions that might improve it. Nevertheless, quantifying workplace performance in an objective yet generalizable manner remains a challenge in the area. For instance, Amazon uses passive sensing on smartphones to measure the performance of their drivers and make sure that they avoid potentially dangerous behaviors like checking the phone while driving. However, phone usage could be the desired behavior in a different job, e.g., consulting.

### v3

@@ Abstract
The modern workplace is undergoing a radical transformation, driven by technological advances that blur the boundaries between human capability and digital augmentation. At the forefront of this evolution is passive sensing technology - a suite of tools that quietly monitor and interpret human behavior without active user engagement. This paper examines how these technologies are reshaping our understanding of workplace dynamics, with a particular focus on employee wellbeing and productivity. Through a comprehensive review of recent research, we explore both the transformative potential and inherent challenges of passive sensing in professional environments. Our analysis reveals emerging patterns in how these technologies can support worker health and performance, while also highlighting critical gaps in current research and opportunities for future innovation. We conclude by outlining a roadmap for integrating passive sensing into future workplaces in ways that enhance human potential while preserving dignity and autonomy.
Keywords:
Passive Sensing Future of Work Workplace Wellbeing Productivity Employee health Wearables Digital behavior AI.
Published at HCI International 2025, DOI: https://doi.org/10.1007/978-3-031-93845-0_29
@@
footnotetext: $\dagger$ These authors contributed equally to this work.
1 Introduction
As technology becomes increasingly integrated into the workplace, researchers studying the “Future of Work" are exploring new ways to enhance work environments and support employee capabilities. Among emerging technologies, passive sensors have shown particular promise for their ability to blend seamlessly into work settings. These sensors gather data without requiring active input from users, offering an unobtrusive way to understand and support workers throughout their day [14].
What makes these sensors particularly valuable is their ability to capture detailed insights into worker behavior while enabling evidence-based analysis grounded in real workplace contexts. Building on this capability, researchers have uncovered strong connections between passive sensing data and key behavioral indicators, including mental wellbeing [91], personality traits [92], and productivity [58]. Traditional performance evaluations often rely on self-reporting or supervisor assessments, which provide only periodic snapshots of worker experience. In contrast, passive sensing technologies can track and understand workers’ experiences continuously over long periods and across different settings, offering a more complete picture of workplace dynamics without disrupting normal work patterns. Given the growing importance of these technologies, we present a comprehensive survey of current research on how passive sensing is being used in workplaces to assess and improve worker wellbeing and productivity. For this study, we define "workplace" as any setting where people perform their job duties, whether in private or shared spaces. This inclusive definition reflects the evolving nature of work environments and ensures our analysis captures the full spectrum of workplace applications.
The significance of this research has been amplified by recent advances in Artificial Intelligence (AI) and Large Language Models (LLM), which have created new opportunities for analyzing and acting on passive sensing data. As organizations adapt to changing workforce needs, combining passive sensing with AI could enable more sophisticated ways to support workers. However, this technological convergence also raises important questions about privacy, ethics, and finding the right balance between technological assistance and human autonomy in tomorrow’s workplace. Through our analysis, we identify these key challenges and opportunities, providing a foundation for understanding how passive sensing technologies can thoughtfully be integrated into the future of work.
@@ 2 Survey Methodology
To ensure the relevance and rigor of our research, we conducted searches for terms like “passive sensing & future of work", “wearables & worker productivity", and “wearables & employee health", leading us to review over 25,000 articles on platforms like Google Scholar. Our selection criteria favored recent papers published since 2015, including some seminal earlier works, and prioritized publications with a Google h5 index of 20 or higher. Furthermore, we imposed specific conditions to refine our focus. Studies had to involve the actual workforce engaged in day-to-day job activities, excluding simulations with non-working groups, such as students in office environments. The technologies considered had to be genuinely passive—unobtrusive, portable, and comfortably integrated into daily routines without requiring active interaction. The technologies should blend with the workers’ routine, rather than the workers changing their everyday routine to satisfy the study requirements. Therefore, research utilizing intrusive devices like EEG headsets, chest straps, gaze trackers, cameras and other cumbersome scientific devices was excluded in favor of those using more discreet and workplace-appropriate technologies. We focus on devices that we expect to have a higher acceptance in the office environment. This stringent selection process allowed us to accurately portray the state-of-the-art developments, progress, and hurdles in this field, providing a foundation to stimulate and guide further research on passive sensing technologies in the workplace.
3 Foundations of Workplace Passive Sensing
Ubiquitous devices, increasingly embedded in everyday life, now capture daily behavior at high resolution through multiple sensors. Researchers have leveraged passive sensing techniques to model users’ behavior and contexts. Even before smartphones and wearables, cell phone sensors (e.g., GPS, Bluetooth) were used for passive context sensing. RealityMining [29] employed location and Bluetooth data to identify social patterns, socially significant places, and organizational rhythms. CenceMe [56] expanded this approach by integrating cameras, microphones, accelerometers, GPS, and other sensors to detect user activities and habits, then shared these on social networks. Beyond personal behavior, passive sensing also enabled large-scale people-centric urban sensing, as MetroSense [30] combined mobile and static sensors throughout a city to support various applications. As mobile and wearable devices advanced, longitudinal tracking became feasible, demonstrated by studies like StudentLife [91] which monitored college students’ behaviors for over two months, assessing mental health, academic performance, and behavioral trends. Passive sensing has since extended into workplaces to monitor physiological factors, improve safety, and enhance efficiency [45]. For example, Tesserae [55] studied over 700 information workers to measure workplace performance, psychological traits, and physical characteristics. In the rest of the article, we focus on applications of passive sensing technologies in the workplace to assess wellbeing and productivity of the workforce as the transition to the Future of Work takes place.
@@ 3 Foundations of Workplace Passive Sensing
Table 1: Studied behavior and sensors used. The following table lists all the behaviors studied in the papers we surveyed and the corresponding sensor data used.
@@
Skin conductance, skin temperature, acceleration [88]; workplace behavioral logs/ computer activity [59, 66, 62], context [86];
@@
EDA, blood volume pulse (BVP), acceleration, phone usage [28]; Computer activity [87, 46, 82]; Speech, calendar [46];
@@
4 Applications of Passive Sensing in the Workplace
@@
In this section, we explore how passive sensing technologies advance our understanding of workplace wellbeing and productivity. Drawing on recent research, we highlight their ability to illuminate the intricate relationship between work environments and employee behavior. This exploration underscores the potential of passive sensing to optimize workplace dynamics and promote a healthier, more productive work culture.
4.1 Understanding Employee Wellbeing
@@
Over the years, there has been an increase in studies that explore the sensing capabilities of smartphones and wearables to assess as well as improve worker’s health and wellbeing. Among the papers selected, mental health issues such as stress, anxiety, affects are important dimensions studied. Most of these studies ask participants to self-report their health and wellbeing using validated instruments. While some studies use one-off survey scores, others employ experience sampling to study worker’s wellbeing in a longitudinal fashion. We list the key studies in Table 1.
@@
Assessing workers’ stress and/or anxiety is perhaps the most studied topic about worker’s wellbeing. As stress and anxiety are known to have an impact on heart rate (HR), most of the studies rely on HR-based signals obtained from wrist-worn wearables that employ photoplethysmography (PPG) [35] sensors. For instance, Feng et al. [34] show that PPG based HR and step counts from a Fitbit can be used to classify the anxiety level of nursing professionals. The authors introduce a pipeline for discovering behavioral consistency features, the inclusion of which leads to an overall improvement in the predictive model, with an accuracy of about 58%. For this work, the authors use the median-splitted score of self-reported State-Trait Anxiety Inventory (STAI) [38] as the ground truth. Similarly, in a study of 10 construction workers, Jebelli et al. [41] use a wrist wearable capable of not only capturing the participant’s HR through PPG signals, but also their electrodermal activity (EDA) and peripheral skin temperature (ST). The study uses salivary cortisol as the ground truth to assess the baseline stress levels of the workers, obtaining an accuracy of about 73% when distinguishing between low, medium and high stress level of the workers. Similarly, Mark et al. [51] discover in their research on information workers that there is a direct correlation between the daily time spent on emails (captured via activity logging software) and increased levels of reported stress. In a bid to craft a stress-reduction intervention for the workplace, Howe et al. [40] conduct a four-week study with 86 participants to evaluate workplace stress-reduction interventions, using signals from email, calendar, facial images, and heart rate data. Their findings show that digital micro-interventions effectively reduce short-term workplace stress. Building on this work, Suh et al.[86] analyze factors influencing intervention engagement, finding that contextual signals and individual traits significantly impact intervention effectiveness, while user-rescheduled interventions show higher engagement rates. Together, these studies demonstrate the value of passive sensing-based interventions while highlighting the importance of individual and contextual factors in their design.
Sleep quality is another frequently studied health issue. Feng et al.[33] study 138 nursing professionals using Fitbit data (PPG-based HR, sleep duration, efficiency, and REM sleep) to predict sleep quality based on PSQI survey[11]. Their motif-based approach achieves 77.47% F1 score, improving upon the 61.90% baseline using Fitbit summary data alone. Martinez et al. [54] examine sleep detection accuracy among 700 information workers, finding that combining phone usage data with wearable data produces better sleep models than using either data source independently. Research has also explored the role of passive sensing in assessing workers’ affect, emotions, and mood. Bin Morshed et al.[8] analyze mood instability in 603 information workers, finding a negative correlation between sleep/activity duration (from Garmin wearables) and mood instability, measured via self-reported PANAS scores[16]. Umematsu et al.[88] use wrist wearables to track skin conductance, temperature, and acceleration in 39 Japanese workers over 30 days. Their predictive model forecasts next-day stress, mood, and health scores with mean absolute errors of 13.47, 14.09, and 18.51, respectively. For fine-grained mood classification, Zenonos et al.[94] develop a framework using physiological data (HR, pulse, PPG, ECG, skin temperature) to classify eight moods into five categories. Mark et al. [49] examine workplace email usage and find a negative correlation with mood balance—higher email engagement corresponds to lower positive affect.
@@
In another study of 50 hospital workers which also uses PANAS, Nadarajan et al. [60] find that speech activity can explain some variance in predicting positive affect measure. Employees wear a specifically designed audio badge during their work-shift hours. The authors extract several features from the audio to identify foreground speech. They then use a linear mixed effects model to estimate positive and negative affect from foreground activation (i.e., the percentage of recording time that foreground speech is present). Similarly, in another study, Robles-Granda et al. [72] utilize multiple sensing modalities to assess anxiety, sleep and affect of 757 information workers. Sensing modalities include wearable, phone application, Bluetooth beacons and social media. Models trained on the fusion of all the features from different sensing modalities lead to up to 13.9% improvement in the symmetric mean absolute percentage error (SMAPE) when predicting affect, anxiety and sleep quality scores. Booth et al. employ multiple sensing modalities to study stress, anxiety and affect of hospital workers [9]. Using Bluetooth tags for location tracking, Fitbit data, and smart shirt sensors, authors achieve 58% accuracy in classifying participants’ mental wellness between high and low stress/affect states. Similarly, Nepal et al. [62] study burnout in cybersecurity incident responders using both passive sensing and self-reports. Analyzing digital activities of 35 security workers, they find that burned-out employees exhibit distinct behavioral patterns - sending more emails after hours, having more weekend work, and showing poorer work-life boundaries.
@@
Other wellbeing-related studies using passive sensing include focus and awakeness. Soto et al.[83] leverage biometric data from an arm-worn device (physical activity, HR, skin response, skin temperature, respiration) to estimate stress, focus, and awakeness in 14 knowledge workers over eight weeks. Their personalized models outperformed generic ones, improving precision, recall, and accuracy by up to 52%. There are also several other works that fuse passive sensing data with some other modality to assess wellbeing. For instance, Saha and Grover et al. [75] propose contextualizing on offline behaviors as obtained from passive sensors to make models better adapted to the social media signals. Authors first cluster participants based on the passive sensing data obtained from their wearable, Bluetooth beacons and smartphone application. Thereafter, they use social media derived features to predict different constructs within each cluster, obtaining an improvement on the baseline generic model by up to 5.43% in predicting self-reported anxiety, affect, and sleep quality of information workers.
@@
4.2 Measuring Workplace Performance